Attendees: Nena Moonier, Bruno Fieramosca, Dean Wyncott, Wendy VanWingeren, Shane Flood, and Patti Pedergnana
- Nena will complete the summer run schedule by the end of this week.
- All FCs will disable their stations tomorrow morning. Patti will take all beamlines global offline about 1000.
- Nena will have a contractor coming out to work on the bio-safety cabinets soon.
- Shane is going to follow up on the WR and CCWP for 26-ID’s shutter bellows adjustments.
- Dean will initiate a WR/CCWP for welding unistrut at 3-ID.
- Property management was looking for a nano-robot in Jin Wang’s office but were unable to locate it.
- The hoisting and rigging assessment is underway. Bruno is checking that slings are up-to-date. He is checking tags and documentation too.
- 6-ID’s Front End CCWP is still not ready. M. Erdmann needs to get an approval memo from J. Cross. The 6-BM-B hutch is opened and the table is turned so that 6-ID FE is accessible.
- 4-ID’s FOE has a lead stack that is going to be moved. The collimator will not be moved, only measured off of the collimator for the bricks location. This CCWP will need to be posted tomorrow.
- All the truck locks have an emergency open/close button in their A hallway. The outer truck lock door location is not visible from the hallway, so this door could be closed on somebody. The close button has been unwired. Only the door open button works now. There are signs posted to depress the door open button in case of an ODH monitor alarm.
- Bruno and A. Stevens walked Sectors 1-8 for the LTHW shutdown.
- There are four ongoing radioactive sample experiments.
- Sectors 20-23 will have switchgear maintenance this weekend.